OpenStudy (anonymous):

Sorry, I put this in the wrong place. Thank You for your great help. Can someone show me step by step instructions. I'm having a problem. Thanks 2/3x-5/6=-1/3

OpenStudy (anonymous):

First, Find a common denominator...so I would have my common denominator be 6

OpenStudy (anonymous):

Once you do that your problem should look like this: 4/6x-5/6=-2/6

OpenStudy (anonymous):

Now your goal is to get the x by itself so first we will add -5/6 to -2/6 to get -7/6

OpenStudy (anonymous):

now your problem should look like this: 4/6x=-7/6 in order to get the x alone you have to divide both sides by 4/6 your final answer should be: -7/4 or -1(3/4)

OpenStudy (anonymous):

Lifesaver indeed. Thank You

OpenStudy (anonymous):

No problem!